OpenStudy (anonymous):

-(3x - 2) = -4(x + 1) - 4 help please!!! :)

OpenStudy (saifoo.khan):

-3x + 2 = -4x -1

OpenStudy (anonymous):

-3x + 2 = -4x - 4 - 4 -3x + 2 = -4x - 8 x + 2 = -8 x = -10

OpenStudy (anonymous):

hick's right, saifoo check your eq

OpenStudy (anonymous):

saifoos screwed it

OpenStudy (anonymous):

thanks guys!